Be sure to answer all parts.

How many mol of platinum (Pt) atoms are equivalent to one billion platinum atoms? Enter your answer in scientific notation.
_×10_mol Pt

(select)−11−19−9−17−13−15
Answer #1

number of atoms = 1.0*10^9

Use:

number of mol = number of atoms / Avogadro’s number

= (1.0*10^9) / (6.022*10^23)

= 1.66*10^-15 mol

Answer: 1.7*10^-15 mol